OpenStudy (anonymous):

solve [0,2pi) square 12 tanx-2=0

OpenStudy (solomonzelman):

can you wirte it out so that I can understand it, you wrote it kind of abstrusely.

OpenStudy (anonymous):

\[\sqrt{12}\tan x - 2 =0\]

OpenStudy (solomonzelman):

so \[\sqrt{12tanx}-2=0\]add two to both sides\[\sqrt{12tanx}=2\]take out 2 from the square root.\[\sqrt{12tanx}=2--->\sqrt{4} \times \sqrt{3tanx}=2--->2\sqrt{3tanx}=2\]divide both sides by 2\[\sqrt{3tanx}=1\]square both sides\[(\sqrt{3tanx})^2=1^2⇒3tanx=1\]divide both sides by 3.\[tanx=1/3\]take inverse tangent of 0.33

OpenStudy (solomonzelman):

I don't have a calculator, sorry.

OpenStudy (solomonzelman):

@shorouq91, is this good?

OpenStudy (anonymous):

i'm working out

OpenStudy (anonymous):

why\[\sqrt{\tan }\]

OpenStudy (solomonzelman):

Oh, my bad!

OpenStudy (solomonzelman):

no it is\[tanx= \frac{1}{2\sqrt{3}}\]so far so good?

OpenStudy (anonymous):

what step i have to used

OpenStudy (solomonzelman):

do you have to show you work?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just give me th \[\sqrt{12}\]

OpenStudy (solomonzelman):

\[\sqrt{12}tanx=2\]simplify the sqrt 12,\[\sqrt{12}tanx=3 ⇒ \sqrt{4} \times \sqrt{3} tanx=2⇒ 2\sqrt{3}tanx=2\]divide both sides by 2, \[\sqrt{3}tanx=1\]

OpenStudy (solomonzelman):

\[\sqrt{3} tanx=1\]divide both sides by sqrt 3

OpenStudy (solomonzelman):

\[tanx=\frac{1}{\sqrt{3}}\]

OpenStudy (solomonzelman):

find what \[\frac{1}{\sqrt{3}}=?\]and take inverse tan of that.

OpenStudy (anonymous):

cos/sin right

OpenStudy (solomonzelman):

The other wAY

OpenStudy (solomonzelman):